OpenStudy (mini):

can U help me plz and fast? Calculate the theoretical probability for spinning black. Calculate the theoretical probability for spining red. P (black) and P (red)

OpenStudy (mini):



OpenStudy (welshfella):

From these figures you can calculate the experimental probability. The best estimate comes with the most spins.

OpenStudy (mini):

the first step is it: P(black) 20/10?

OpenStudy (mini):

ok

OpenStudy (welshfella):

No first one is 20 / total spins = 20/30

OpenStudy (mini):

ok

OpenStudy (mini):

then I need to answer 20 /30?

OpenStudy (mini):

welshfella, what do I need to do next?

OpenStudy (welshfella):

the closest value to the theorretical probability will be from Olga's 300 spins P(Black) = 174/300 and P(red) = 126/300 as a decimal these are 0.58 and 0.42

OpenStudy (mini):

So wait what are we doing with our 20/30?

OpenStudy (welshfella):

If its a fair spinner with equal areas for black and white the theoretical probability is 0.50 for black and red.

OpenStudy (welshfella):

Just state the 3 results and write that Olga's result is closer to the theoretical probability because of the large number of spins involved.

OpenStudy (mini):

No I can't because I need to show my work :) how I did it :)

OpenStudy (mini):

So, 20/30

OpenStudy (mrnood):

@welshfella I answered a very similar question from a fairly rude poster about an hour ago - he seemed to have a couple of similar accounts, and used similar phrases and table layout. Please make sure you are not being led astray here

OpenStudy (mini):

Then we need to simplify right?

OpenStudy (welshfella):

for first one P = 2/3 and 1/3 as i've said second P = 38/60 and 22/ 60 and for Olga's work i've mentioned that already

OpenStudy (welshfella):

OK thanks @MrNood

OpenStudy (mini):

It doesn't make sence with the 38, how come there is a 38?

OpenStudy (mini):

It shouldn't be Liam first? or something like that?

OpenStudy (welshfella):

Maria spins 38 Black's from 60 spins here P(black) = 38 / total spins = 38 / 60

OpenStudy (mini):

And Liam?

OpenStudy (welshfella):

check my posts - ive shown you how to work out liams and olga;s.

OpenStudy (mini):

We need to simplify right the 20/30?

OpenStudy (mini):

ok

OpenStudy (mini):

So I have P(black) = 20/30

OpenStudy (aaronandyson):

Probability of an event = No.of favorable outcomes/total number of possible outcome

OpenStudy (mini):

Ok, that I understand I think, but the thing about 20/30 we need to simplify it right?

OpenStudy (aaronandyson):

Yes,that 2/3

OpenStudy (mini):

Ok, now I need to do 2/3 = 0.66666 but It will be 0.6 right?

OpenStudy (aaronandyson):

No.In probability we keep the fractions.

OpenStudy (mini):

Ok, but if we don't it would be o.6 still?

OpenStudy (mini):

Then the P(Red) = 10/30?

OpenStudy (aaronandyson):

Thats 1/3

OpenStudy (mini):

Ok, thank you :)

OpenStudy (mini):

Is that all?

OpenStudy (aaronandyson):

2/3 and 1/3 for liam. we still have 2 more people.

OpenStudy (mini):

No it's alright it's these too that I needed the most :)

OpenStudy (mini):

do you want a medal?

OpenStudy (aaronandyson):

Yes.

OpenStudy (mini):

thank you, you save my life;)

OpenStudy (aaronandyson):

:)